Using the multiplier model, evaluate the effects of the following scenarios on Real GDP. In all of these scenarios, the MPC is 0.8.
A) The government increases its spending on infrastructure by $100 million. B) The government increases taxes by $100 million.
C) The government reduces its spending on goods and services by $10 million. D) The government reduces social security benefits by $10 million.
1
Expert's answer
2019-03-05T13:41:19-0500

GDP = C + I + G

where

C- consumption,

I – investment,

G – government spending.

MPC is marginal propensity to consume.

A) The investment multiplier will be:

1/(1-MPC)=1/(1-0.8)=5

Since the initial increase in spending is $100 million and the multiplier is 5, this is simply:

∆GDP=5*100=$500

B) The Tax Multiplier is:

MPC/MPS=MPC/(1-MPC)=0.8/0.2=4

∆GDP=4*100=$400

C) Since spending reduced by $10 million and the multiplier is 5, then:

∆GDP=5*(-10)=-$50

D) Social Security payments are the part of citizens’ income. Eventually they will be consumed, but not completely. The total consumption will be reduced by: $10*0.8=$8 million.

The change of real GDP will be:

∆GDP=5*(-8)=-$40